# Avoiding the Price of Adaptivity: Inference in Linear Contextual Bandits via Stability

### v1

@@ Abstract
Statistical inference in contextual bandits is complicated by the adaptive, non-i.i.d. nature of the data. A growing body of work has shown that classical least-squares inference may fail under adaptive sampling, and that constructing valid confidence intervals for linear functionals of the model parameter typically requires paying an unavoidable inflation of order $\sqrt{d\log T}$. This phenomenon—often referred to as the price of adaptivity—highlights the inherent difficulty of reliable inference under general contextual bandit policies.
@@
In this paper, we propose and analyze a penalized EXP4 algorithm for linear contextual bandits. Our first main result shows that this procedure satisfies the Lai–Wei stability condition and therefore admits valid Wald-type confidence intervals for linear functionals. Our second result establishes that the same algorithm achieves regret guarantees that are minimax optimal up to logarithmic factors, demonstrating that stability and statistical efficiency can coexist within a single contextual bandit method. Finally, we complement our theory with simulations illustrating the empirical normality of the resulting estimators and the sharpness of the corresponding confidence intervals.
@@ 1 Introduction
where $\phi:\mathcal{A}\times\mathcal{X}\to\mathbb{R}^{d}$ is a known feature map and $\beta^{\star}\in\mathbb{R}^{d}$ is an unknown parameter to be learned. We assume that the noise sequence $\{\varepsilon_{t}\}_{t\geq 1}$ satisfies $\mathbb{E}[\varepsilon_{t}\mid\mathcal{F}_{t-1},x_{t}]=0$. Our goal is to construct confidence intervals for linear functional of the form $a^{\top}\beta^{\star}$, for any fixed vector $a\in^{d}$.
@@ 1.1 Price of Adaptivity:
[…]
@@ 1.3 Contributions
In this paper, we study the problem of constructing confidence intervals for linear functionals of the model parameters in a linear contextual bandit problem. Section 2 provides a detailed description of the contextual bandit framework we consider, while Section 3 introduces the algorithm we analyze. The procedure, stated formally in Algorithm 1, is a penalized variant of the classical EXP4 algorithm designed to enforce stability while allowing adaptive exploration. Our first main result, stated in Theorem 1, establishes that the penalized EXP4 algorithm is stable in the sense of lai1982least (see definition (1)), enabling the construction of asymptotically valid Wald-type confidence intervals. Our second result, detailed in Theorem 2 , analyzes the regret of the proposed algorithm. We show that, despite being explicitly designed for inferential stability, the procedure simultaneously achieves minimax–optimal regret guarantees up to logarithmic factors. Thus, stability and statistical efficiency can coexist within a single contextual bandit algorithm. Section 6 contains proofs of our main theorems with some technical details deferred to the Appendix. Finally, in Section 5, we complement our theoretical findings with numerical experiments demonstrating the empirical validity of the Wald confidence intervals produced by Algorithm 1.
@@ 1.5 Notation
For any matrix $A$, let $\|A\|_{\mathrm{op}}$ and $\|A\|_{F}$ denote the operator norm and the Frobenius norm, respectively. For any vector $v\in\mathbb{R}^{d}$, the standard $\ell_{p}$-norm for $p\in(0,\infty)$ is denoted by $\|v\|_{p}$. For a real-valued random variable $u$, we define $\|u\|_{p}:=\mathbb{E}\!\left[|u|^{p}\right]^{1/p}.$ Given a fixed weight vector $w\in\mathbb{R}^{d}$, we define a weighted norm on $\mathbb{R}^{d}$, denoted by $\|\cdot\|_{w,\ast}$, as $\|v\|_{w,\ast}^{2}:=\sum_{i=1}^{d}w_{i}v_{i}^{2},$ where $v\in\mathbb{R}^{d}.$ For two nonnegative sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $b_{n}\gg a_{n}$ if $\frac{b_{n}}{a_{n}}\to\infty\quad\text{as }n\to\infty.$ Throughout, we suppress absolute constants in inequalities by using the notation $\lesssim$ and $\gtrsim$. The Loewner partial order on symmetric matrices is denoted by $\preceq$.
@@ 2 Problem Setup
where $c(x_{t},a_{t})\in\mathbb{R}^{d}$ is a known feature representation of the pair $(x_{t},a_{t})$, and $\beta^{\star}\in\mathbb{R}^{d}$ is an unknown parameter vector. We use $\mathcal{F}_{t}:=\sigma(x_{1},a_{1},\ell_{1},\ldots,x_{t},a_{t},\ell_{t})$ to denote the $\sigma$-field generated by observations up to time $t$. We assume that the noise sequence $\{\varepsilon_{t}\}_{t=1}^{T}$ satisfies
$\mathbb{E}[\varepsilon_{t}\mid\mathcal{F}_{t-1},x_{t}]=0.$ (9)
At each round, the learner chooses an action by sampling from a mixture of $K$ base experts $\{\pi_{1},\ldots,\pi_{K}\}$, where each $\pi_{k}(\cdot\mid x_{t})$ defines an arbitrary (possibly stochastic) policy mapping the context $x_{t}$ to a distribution over actions. The learner maintains mixture weights $w_{t}=(w_{t,1},\ldots,w_{t,K})$ belonging to the $\varepsilon$-simplex
@@
We take $\epsilon$ to be a small positive tuning parameter that decays with the number of rounds $T$. Introduction of the tuning parameter $\varepsilon$ guarantees $w_{k}\geq\varepsilon$ and consequently ensures that various importance ratio based estimators used in our algorithm are always well-defined.
The effective policy at round $t$ is therefore the convex mixture
@@
The learner then draws an action $a_{t}\sim Q_{t}(\cdot\mid x_{t})$ and observes the corresponding stochastic loss $\ell_{t}$. Throughout, we work under the following assumptions:
@@
    $\mathbb{E}[\varepsilon_{t}|\mathcal{F}_{t-1},x_{t}]=0$, $|\varepsilon_{t}|\leq 1$, and $\mathbb{E}[\varepsilon^{2}_{t}|\mathcal{F}_{t-1},x_{t},a_{t}]=\sigma^{2}$ for all $t\geq 1$.
@@
Assumption (A1) states that the noise sequence is conditionally zero mean, bounded and homoscedastic. Assumption (A2) imposes a boundedness assumption on the parameter vector $\beta^{\star}$ and feature map $c(\cdot,\cdot)$. Assumptions (A1) and (A2) together ensure that the observed loss $\{\ell_{t}\}$ is uniformly bounded. Assumption (A3) posots that the the context vectors $\{x_{t}\}$ are i.i.d., and that the weighted covariance matrix associated with each expert is non-singular. Assumption (A4) can always be satisfied by taking one of the expert is as a uniform expert $\pi_{\mathrm{unif}}(a\mid x)=1/|\mathcal{A}|$ for all action $a\in\mathcal{A},\;\;\text{and context}\;\;x\in\mathcal{X}$. This assumption along with our choice of weights $w_{tk}\geq\epsilon>$ (see equation (10)) ensures that $Q_{t}(a_{t}\mid x_{t})>0$ for all $t\geq 1$ and various importance weight based estimators are well-defined.
@@ 3 A Penalized Exp4 Algorithm
Algorithm 1 Penalized EXP4
1:Input: Base policies $\{\pi_{k}\}_{k=1}^{K}$, stepsizes $\{\eta\}$, penalty $\lambda>0$, floor $\varepsilon\in(0,1/K]$.
@@ 4 Main results
In this section we state our main results. Section 1.1 highlights that, for general adaptive data-collection rules, valid confidence intervals for linear functionals must inflate by a factor on the order of $\sqrt{d\log T}$. In this section, we show that the penalized EXP4 procedure (Algorithm 1) avoids this worst-case behavior. In particular, this algorithm satisfies the stability from Definition 1, which ensures the validity of Wald-type inference, and it does so while retaining near–minimax-optimal regret.
@@ 4.1 Stability and a central limit theorem
Under Assumptions (A1)-(A4), the penalized-EXP4 algorithm (1) with step size $\eta=\sqrt{\frac{\log K}{|\mathcal{A}|T}}$, and tuning parameters $\epsilon=\frac{1}{KT}$, $\lambda=\frac{\gamma_{T}}{\sqrt{T}}$ with $\gamma_{T}\rightarrow\infty$ and $T/\log^{2}T\gg\gamma_{T}$, is stable; we have
@@
with $\Sigma_{T}^{\star}$ from (25). Consequently, for any fixed $a\in^{d}$ we have
$\frac{a^{\top}(\widehat{\beta}_{\mathrm{OLS}}-\beta^{\star})}{\widehat{\sigma}\sqrt{a^{\top}S_{T}^{-1}a}}\xrightarrow{d}\mathcal{N}\left(0,1\right)$ (26a)
where $\widehat{\beta}_{\mathrm{OLS}}$ is the least square estimate, and $\widehat{\sigma}$ is any consistent estimator of noise standard deviation $\sigma$.
Throughout we assume that the sample covariance matrix $S_{T}$ is invertible. If $S_{T}$ is not invertible, one can construct a ridge estimator with the same guarantees. We highlight the technical details in Section 6. A direct consequence of Theorem 1 is that the Wald confidence interval (4) is asymptotically exact: for any fixed $a\in\mathbb{R}^{d}$,
@@ 4.2 Regret guarantees
We prove Theorem 2 in Section 6.2. Setting $\gamma_{T}=\sqrt{\log T}$ yields regret of order $O\!\left(\sqrt{TK\log(KT)}\right)$, matching minimax lower bounds up to logarithmic factors. Thus, despite being explicitly regularized for inferential stability, the procedure retains the characteristic efficiency of the EXP4 family.

### v2

@@ Abstract
Statistical inference in contextual bandits is complicated due to the adaptive, non-i.i.d. nature of the data. A growing body of work has shown that classical least-squares inference may fail under adaptive sampling, and that constructing valid confidence intervals for linear functionals of the model parameter typically requires paying an unavoidable inflation of order $\sqrt{d\log T}$. This phenomenon—often referred to as the price of adaptivity—highlights the inherent difficulty of reliable inference under general contextual bandit policies.
@@
In this paper, we propose and analyze a regularized EXP4 algorithm for linear contextual bandits. Our first main result shows that this procedure satisfies the Lai–Wei stability condition and therefore admits valid Wald-type confidence intervals for linear functionals. We additionally provide quantitative rates of convergence in the associated central limit theorem. Our second result establishes that the same algorithm achieves regret guarantees that are minimax optimal up to logarithmic factors, demonstrating that stability and statistical efficiency can coexist within a single contextual bandit method. As an application of our theory, we show how it can be used to construct confidence intervals for the conditional average treatment effect (CATE) under adaptively collected data. Finally, we complement our theory with simulations illustrating the empirical normality of the resulting estimators and the sharpness of the corresponding confidence intervals.
@@ 1 Introduction
where $\phi:\mathcal{A}\times\mathcal{X}\to\mathbb{R}^{d}$ is a known feature map and $\beta^{\star}\in\mathbb{R}^{d}$ is an unknown parameter to be learned. We assume that the noise sequence $\{\varepsilon_{t}\}_{t\geq 1}$ satisfies $\mathbb{E}[\varepsilon_{t}\mid\mathcal{F}_{t-1},x_{t},a_{t}]=0$. Our goal is to construct confidence intervals for linear functional of the form $a^{\top}\beta^{\star}$, for any fixed vector $a\in^{d}$.
@@ 1.1 Price of Adaptivity:
where $\widehat{\beta}_{\mathrm{OLS}}$ is the least square estimator, $z_{1-\alpha/2}$ is the standard normal quantile and $\hat{\sigma}$ is a consistent estimator of noise variance $\sigma$. Comparing the length of the two confidence intervals from (2) and (4), we observe  [^1] One usually takes $\lambda$ to be small and the effect of the term $\lambda\sqrt{S}$ in (3) is negligible.
@@ 1.3 Contributions
In this paper, we study the problem of constructing confidence intervals for linear functionals of the model parameters in a linear contextual bandit problem. Section 2 provides a detailed description of the contextual bandit framework we consider, while Section 3 introduces the algorithm we analyze. The procedure, stated formally in Algorithm 1, is a regularized variant of the classical EXP4 algorithm designed to enforce stability while allowing adaptive exploration. Our first main result, stated in Theorem 1, establishes that the regularized EXP4 algorithm is stable in the sense of lai1982least (see definition (1)), enabling the construction of asymptotically valid Wald-type confidence intervals. In addition, we provide a non-asymptotic rates of convergence to normality. Our second result, detailed in Theorem 2 , analyzes the regret of the proposed algorithm. We show that, despite being explicitly designed for inferential stability, the procedure simultaneously achieves minimax–optimal regret guarantees up to logarithmic factors. Thus, stability and statistical efficiency can coexist within a single contextual bandit algorithm. As an application, we establish asymptotically exact Wald-type confidence intervals for conditional average treatment effects in linear models with adaptively assigned treatments, using ridge regression. Section 6 contains proofs of our main theorems with some technical details deferred to the Appendix. Finally, in Section 5, we complement our theoretical findings with numerical experiments demonstrating the empirical validity of the Wald confidence intervals produced by Algorithm 1.
@@ 1.5 Notation
For any matrix $A$, let $\|A\|_{\mathrm{op}}$ and $\|A\|_{F}$ denote the operator norm and the Frobenius norm, respectively. For any vector $v\in\mathbb{R}^{d}$, the standard $\ell_{p}$-norm for $p\in(0,\infty)$ is denoted by $\|v\|_{p}$. For a real-valued random variable $u$, we define $\|u\|_{p}:=\mathbb{E}\!\left[|u|^{p}\right]^{1/p}.$ Given a fixed weight vector $w\in\mathbb{R}^{d}$, we define a weighted norm on $\mathbb{R}^{d}$, denoted by $\|\cdot\|_{w,\ast}$, as $\|v\|_{w,\ast}^{2}:=\sum_{i=1}^{d}w_{i}v_{i}^{2},$ where $v\in\mathbb{R}^{d}.$ For two nonnegative sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $b_{n}\gg a_{n}$ if $\frac{b_{n}}{a_{n}}\to\infty\quad\text{as }n\to\infty.$ Throughout, we suppress absolute constants in inequalities by using the notation $\lesssim$ and $\gtrsim$. The Loewner partial order on symmetric matrices is denoted by $\preceq$. Finally, the Kolmogorov distance between two real-valued random variables $X$ and $Y$ is denoted by $\mathrm{d}_{\mathrm{K}}(X,Y)$ and is defined as
$\mathrm{d}_{\mathrm{K}}(X,Y):=\sup_{t\in\mathbb{R}}\left|\mathbb{P}(X\leq t)-\mathbb{P}(Y\leq t)\right|.$
@@ 2 Problem Setup
where $c(x_{t},a_{t})\in\mathbb{R}^{d}$ is a known feature representation of the pair $(x_{t},a_{t})$, and $\beta^{\star}\in\mathbb{R}^{d}$ is an unknown parameter vector. At each round, the learner chooses an action by sampling from a mixture of $K$ base experts $\{\pi_{1},\ldots,\pi_{K}\}$, where each $\pi_{k}(\cdot\mid x_{t})$ defines an arbitrary (possibly stochastic) policy mapping the context $x_{t}$ to a distribution over actions. The learner maintains mixture weights $w_{t}=(w_{t,1},\ldots,w_{t,K})$ belonging to the $\varepsilon$-simplex
@@
We take $\epsilon$ to be a small positive tuning parameter that decays with the number of rounds $T$. Introduction of the tuning parameter $\varepsilon$ guarantees $w_{k}\geq\varepsilon$ and consequently ensures that various importance ratio based estimators used in our algorithm are always well-defined. The effective policy at round $t$ is therefore the convex mixture
@@
The learner then draws an action $a_{t}\sim Q_{t}(\cdot\mid x_{t})$ and observes the corresponding stochastic loss $\ell_{t}$. We use $\mathcal{F}_{t}:=\sigma(x_{1},a_{1},\ell_{1},\ldots,x_{t},a_{t},\ell_{t})$ to denote the $\sigma$-field generated by observations up to time $t$.
Throughout, we work under the following assumptions:
@@
    $\mathbb{E}[\varepsilon_{t}|\mathcal{F}_{t-1},x_{t},a_{t}]=0$, $|\varepsilon_{t}|\leq 1$, and $\mathbb{E}[\varepsilon^{2}_{t}|\mathcal{F}_{t-1},x_{t},a_{t}]=\sigma^{2}$ for all $t\geq 1$.
@@
Assumption (A1) states that the noise sequence is conditionally zero mean, bounded and homoscedastic. Assumption (A2) imposes a boundedness assumption on the parameter vector $\beta^{\star}$ and feature map $c(\cdot,\cdot)$. Assumptions (A1) and (A2) together imply that the observed losses $\{\ell_{t}\}$ are uniformly bounded. Consequently, in Assumption (A3) we may assume without loss of generality that $\ell_{t}\geq 0$. Indeed, if the losses are not necessarily nonnegative, uniform boundedness guarantees the existence of a constant $C>0$ such that $\ell_{t}+C\geq 0$ almost surely for all $t$. Replacing $\ell_{t}$ by $\ell_{t}+C$ amounts to adding a constant offset to the loss model in (8), which can be equivalently absorbed into a redefinition of the feature map and parameter vector. This transformation leaves the regret, stability properties, and inferential guarantees established in this paper unchanged.
Assumption (A4) posits that the the context vectors $\{x_{t}\}$ are i.i.d., and that the weighted covariance matrix associated with each expert is non-singular. Assumption (A5) can always be satisfied by taking one of the expert is as a uniform expert $\pi_{\mathrm{unif}}(a\mid x)=1/|\mathcal{A}|$ for all action $a\in\mathcal{A},\;\;\text{and context}\;\;x\in\mathcal{X}$. This assumption along with our choice of weights $w_{t,k}\geq\epsilon>0$ (see equation (9)) ensures that $Q_{t}(a_{t}\mid x_{t})>0$ for all $t\geq 1$ and various importance weight based estimators are well-defined.
@@ 3 A regularized Exp4 Algorithm
Algorithm 1 Regularized EXP4
1:Input: Base policies $\{\pi_{k}\}_{k=1}^{K}$, stepsizes $\{\eta\}$, penalty $\lambda>0$, floor $\epsilon\in(0,1/K]$.
@@ 4 Main results
In this section we state our main results. Section 1.1 highlights that, for general adaptive data-collection rules, valid confidence intervals for linear functionals must inflate by a factor on the order of $\sqrt{d\log T}$. In this section, we show that the regularized EXP4 procedure (Algorithm 1) avoids this worst-case behavior. In particular, this algorithm satisfies the stability from Definition 1, which ensures the validity of Wald-type inference, and it does so while retaining near–minimax-optimal regret.
@@ 4.1 Stability and a central limit theorem
Under Assumptions (A1)-(A5), the regularized-EXP4 algorithm (1) with step size $\eta=\sqrt{\frac{\log K}{|\mathcal{A}|T}}$, and tuning parameters $\epsilon=\frac{1}{KT}$, $\lambda=\frac{\gamma_{T}}{\sqrt{T}}$ with $\gamma_{T}\rightarrow\infty$ and $T/\log^{2}T\gg\gamma_{T}$, is stable; we have
@@
with $\Sigma_{T}^{\star}$ from (24). Furthermore, for any fixed $a\in^{d}$ we have
$\sup_{a\in^{d}/\{0\}}\operatorname{d}_{\operatorname{K}}\!\left(\frac{a^{\top}(\widehat{\beta}_{\mathrm{OLS}}-\beta^{\star})}{\sigma\sqrt{a^{\top}S_{T}^{-1}a}},\;Z\right)\;\leq\;C\left[\Psi(\gamma_{T})^{1/3}+\frac{1}{T^{1/3}}+d\cdot\operatorname{exp}\left\{-\frac{(\lambda^{\star}_{L})^{2}}{32+8\lambda^{\star}_{L}/3}\cdot T\right\}\right]$ (25a)
where $Z\sim\mathcal{N}(0,1)$, $C$ is a constant independent of $T$, and
$\Psi(\gamma_{T}):=\sqrt{10{\frac{\sqrt{|\mathcal{A}|\log K}}{\gamma_{T}}}+\frac{8\gamma_{T}\log(K)\log^{2}\left(KT\right)}{|\mathcal{A}|^{2}T}}$ (25b)
A direct consequence of Theorem 1 is that the Wald confidence interval (4) is asymptotically exact: for any fixed $a\in\mathbb{R}^{d}$,
@@
We now show that the stability and central limit theorem established in Theorem 1 continue to hold when the ordinary least-squares estimator is replaced by a ridge estimator. Introducing a ridge penalty is natural in adaptive experimentation settings, where the sample covariance matrix may be ill-conditioned or singular in finite samples due to exploration constraints. Let $\lambda_{rid}>0$ denote a regularization parameter and define the regularized Gram matrix $R_{T}:=S_{T}+\lambda_{rid}\,\mathbf{I}$. We consider the ridge estimator
$\widehat{\beta}_{\mathrm{rid}}:=R_{T}^{-1}S_{T}\beta^{\star}+R_{T}^{-1}\sum_{t=1}^{T}z_{t}\varepsilon_{t},$ (26)
where $z_{t}:=c(x_{t},a_{t})$. We show that, under the same stability conditions as in Theorem 1, Wald-type inference based on the ridge estimator remains asymptotically valid:
Corollary 1.
Under the set up of Theorem 1, the ridge-estimator $\hat{\beta}_{rid}$ with $\lambda_{rid}\ll\sqrt{T}$ satisfies
$\sup_{a\in^{d}/\{0\}}\operatorname{d}_{\operatorname{K}}\!\left(\frac{a^{\top}(\hat{\beta}_{rid}-\beta^{\star})}{\sigma\sqrt{a^{\top}R_{T}^{-1}a}},\;Z\right)\;\leq\;C\left[\Psi(\gamma_{T})^{1/3}+\operatorname{exp}\left\{-\frac{(\lambda^{\star}_{L})^{2}}{32+8\lambda^{\star}_{L}/3}\cdot T\right\}+\frac{1}{T^{1/3}}+\frac{\lambda_{rid}}{\ \sqrt{\Psi(\gamma_{T})T}}\right]$
We prove this corollary in Appendix B. Corollary 1 has important implications in estimation of heterogeneous treatment effects which we discuss in the next subsection.
4.2 Inference for Conditional Average Treatment Effects
The conditional average treatment effect (CATE) plays a central role in modern causal inference, as it characterizes heterogeneity in treatment responses across contexts. Formally, for a binary treatment $T\in\{0,1\}$ and covariates $X$, the CATE is defined as
$\tau(x):=\mathbb{E}\!\left[W(1)-W(0)\mid X=x\right],$
where $W(a)$ denotes the potential outcome under treatment $a$. CATEs are foundational to applications such as personalized medicine (powers2018some), targeted advertising (ascarza2018retention) and experimentation platforms (green2012modeling), where decisions are tailored to individual or contextual characteristics.
Despite its importance, valid pointwise inference for CATE for $\tau(x)$ is non-trivial even in non-adaptive settings, as CATE is a non-regular estimand in general (athey2016recursive; wager2018estimation; chernozhukov2018double; kunzel2019metalearners; nie2021quasi). Furthermore, standard results do not generalize well to adaptive data-collection regimes—such as contextual bandits or reinforcement learning—where treatment assignments depend on past observations.
Our results provide a principled resolution to this inference problem for a practically relevant class of models. Specifically, when the conditional mean outcome is linear in known features,
$\mathbb{E}[W_{t}\mid X_{t}=x,a_{t}=a]=c(x,a)^{\top}\beta^{\star},$
the CATE admits the representation
$\tau(x)=\big(c(x,1)-c(x,0)\big)^{\top}\beta^{\star},$
that is, the CATE is a linear functional of the unknown parameter $\beta^{\star}$. In this case, inference for CATE reduces to inference for a linear contrast of $\beta^{\star}$.
When treatment assignments are generated adaptively using the regularized EXP4 algorithm (Algorithm 1), the stability guarantees established in Theorem 1 and Corollary 1 ensure that ridge regression retains a central limit theorem with the same asymptotic variance as in non-adaptive designs. This allows for asymptotically valid Wald-type confidence intervals for CATE under adaptively collected data, without paying any additional price for adaptivity abbasi2011improved. We formalize this implication below.
Corollary 2.
Consider the setup of Section 4.2 with noise satisfying Assumptions (A1)(A5). Let treatment assignments be generated by Algorithm 1, and let $\widehat{\beta}_{\mathrm{ridge}}$ denote the ridge estimator with penalty $\lambda_{rid}\ll\sqrt{T}$ as in Corollary 1. Then, for any fixed $x$ and contrast vector a(x):=c(x,1)c(x,0)a(x):=c(x,1)-c(x,0),
$\operatorname{d}_{\operatorname{K}}\!\left(\frac{a^{\top}(\hat{\beta}_{rid}-\beta^{\star})}{\sigma\sqrt{a^{\top}R_{T}^{-1}a}},\;Z\right)\;\leq\;C\left[\Psi(\gamma_{T})^{1/3}+\operatorname{exp}\left\{-\frac{(\lambda^{\star}_{L})^{2}}{32+8\lambda^{\star}_{L}/3}\cdot T\right\}+\frac{1}{T^{1/3}}+\frac{\lambda_{rid}}{\ \sqrt{\Psi(\gamma_{T})T}}\right]$
where $R_{T}:=S_{T}+\lambda_{rid}\cdot\mathbf{I}$.
@@ 4.3 Regret guarantees
We prove Theorem 2 in Section 6.2. Setting $\gamma_{T}=\sqrt{\log T}$ yields regret of order $O\!\left(\sqrt{TK\log(KT)}\right)$, matching minimax lower bounds up to logarithmic factors. The same choice balances the convergence rate in Theorem 1, leading to a dimension-free Wald-type interval with logarithmic convergence. Thus, despite being explicitly regularized for inferential stability, the procedure retains the characteristic efficiency of the EXP4 family.